25cm3 of a solution of potassium carbonate required 26.8cm3 of 0.1M HCL for neutralization. If 50cm3 of K2CO3 are mixed with 20cm3 of 0.5m hcl ,how many cubic cm of 0.25M NAOH solution must be added to make the solution neutral
1
Expert's answer
2020-04-18T07:00:14-0400

First calculate the strength of the potassium carbonate solution

"25\times M=26.8\times0.1"

"M=0.1072"


in second case

millimoles of (potassium carbonate + NaOH )= millimoles of HCL

"0.1072\times50+0.25\times V=20\times0.5
\\V=18.56 cm^3"